OpenStudy (anonymous):

find the integers,given the following sum is-15 product is 50

OpenStudy (anonymous):

x+y=-15 xy=50 x=-10 y=-15

OpenStudy (anonymous):

y=-5

OpenStudy (anonymous):

sorry..hehehehhe

OpenStudy (anonymous):

x=-10 y=-5

OpenStudy (anonymous):

\[x+y=-15, xy=50. hence, x-y=\sqrt{(x+y)^{2}-4xy}=5. then you will get the terms as -10 and -5\]

OpenStudy (anonymous):

you could not get it?

OpenStudy (anonymous):

ok, if you can't get it, at least give me a medal, please. i assure you, i answered the question correct

OpenStudy (anonymous):

ok!!!!.. ^^

OpenStudy (anonymous):

eq1. x + y= -15 eq2. xy = 50 manipulate eq1 x=-15-y (eq3) substitute eq 3 to eq2 (-15-y)y=50 \[-15y-y^{2}=50\] transpose \[50+15y+y^2=0\] \[y ^{2}+15y+50=0\] (x+10)(x+5) x+10=0 x=-10 x+5=0 x=-5

OpenStudy (anonymous):

Oh! Frightening!!!!!!!

OpenStudy (anonymous):

sorry..in the last part...its not x, its y....sorry...hehehehhe

OpenStudy (anonymous):

that's alright. assume that x=y and y=x :)